# Formatting Instructions For NeurIPS 2024

### v1

@@ Abstract
Graph neural networks (GNNs) have emerged as powerful tools for processing relational data in applications. However, GNNs suffer from the problem of oversmoothing, the property that the features of all nodes exponentially converge to the same vector over layers, prohibiting the design of deep GNNs. In this work we study oversmoothing in graph convolutional networks (GCNs) by using their Gaussian process (GP) equivalence in the limit of infinitely many hidden features. By generalizing methods from conventional deep neural networks (DNNs), we can describe the distribution of features at the output layer of deep GCNs in terms of a GP: as expected, we find that typical parameter choices from the literature lead to oversmoothing. The theory, however, allows us to identify a new, non-oversmoothing phase: if the initial weights of the network have sufficiently large variance, GCNs do not oversmooth, and node features remain informative even at large depth. We demonstrate the validity of this prediction in finite-size GCNs by training a linear classifier on their output. Moreover, using the linearization of the GCN GP, we generalize the concept of propagation depth of information from DNNs to GCNs. This propagation depth diverges at the transition between the oversmoothing and non-oversmoothing phase. We test the predictions of our approach and find good agreement with finite-size GCNs. Initializing GCNs near the transition to the non-oversmoothing phase, we obtain networks which are both deep and expressive.
@@ 1 Introduction
This approach allows us to extend the concept of a regular and a chaotic phase to GCNs: in the regular phase, which describes most of the GCNs studied in the current literature, distances between node features shrink over layers and exponentially attain the same value. We therefore call this the oversmoothing phase. On the other hand, if one increases the variance of the weights at initialization, it is possible to transition into the chaotic phase. In this phase, distances between node features converge to a fixed but finite distance at infinite depth. The convergence point is fully determined by the underlying graph structure and the hyperparameters of the GCN and may differ for different pairs of nodes. GCNs initialized in this phase thus do not suffer from oversmoothing. We find that the convergence point is informative about the topology of the underlying graph and may be used for node classification with GCNs of more than $1,000$ layers. Near the transition point, GCNs at large depth offer a trade-off between feature information and information contained in the neighborhood relation of the graph. We test the predictions of this theory and find good agreement in comparison to finite-size GCNs applied to the contextual stochastic block model [8]. Our approach applies to graphs with arbitrary topologies, depths, and nonlinearities.
@@ 2 Related Work
Oversmoothing is a well-known challenge within the GNN literature [19, 32]. On the theoretical side, rigorous techniques have been used to prove that oversmoothing is inevitable. The authors in [27] show that GCNs with the ReLU non-linearity exponentially lose expressive power and only carry information of the node degree in the infinite layer limit. These results have been extended in [2] to handle non-linearities different from ReLU. Also graph attention networks have been proven to oversmooth inevitably [40]. These proofs, however, make assumptions on the weight matrices which, as we will show, exclude networks in the chaotic, non-oversmoothing phase.
@@ 3.1 Network architecture
with $\phi$ an elementwise nonlinear activation function, $\boldsymbol{b}^{(l)}$ an optional bias term, weight matrices $\boldsymbol{W}^{(l)}\in\mathbb{R}^{d_{l}\times d_{l-1}}$, and $1\in\mathbb{R}^{N}$ a vector of all ones. We note that many GNN architectures studied in the literature are unbiased, which can be recovered by setting $\boldsymbol{b}^{(l)}$ to zero. We use a noisy linear readout, so that the output of the network is given by
$Y=A\boldsymbol{X}^{(L)}\boldsymbol{W}^{(L+1)\top}+1\boldsymbol{b}^{(L+1)\top}+\varepsilon\,,$ (2)
with $\varepsilon\in\mathbb{R}^{N}$ a vector of independent Gaussian random variables: $\epsilon_{\alpha}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma_{ro}^{2})$. We set $d_{L+1}=1$ so that the GCN output at each node is scalar; the architecture, however, can easily be extended to multi-dimensional outputs [17].
@@
$y_{\alpha}$ $=h_{\alpha}^{(L+1)}+\varepsilon_{\alpha}\,,$ (5)
where $\boldsymbol{x}_{\alpha}^{(l)}\in\mathbb{R}^{d_{l}}$ is the feature vector of node $\alpha$ in layer $l$ and $y_{\alpha}\in\mathbb{R}$ the network output for node $\alpha$. The values $\boldsymbol{h}_{\alpha}^{(l)}\in\mathbb{R}^{d_{l}}$ are linear functions of the features $\boldsymbol{x}_{\beta}^{(l-1)}$ and represent the input to the activation functions; we therefore refer to them as preactivations. The nonlinearity $\phi(x)$ is applied elementwise to the preactivations $\boldsymbol{h}_{\alpha}^{(l)}$. While we leave $\phi(x)$ general for the development of the theory, we will use $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ in our experiments in Section 4. The scaling factor in the $\mathrm{erf}$ is chosen such that $\frac{\partial\phi}{\partial x}(0)=1$. We use independent and identical Gaussian priors for all weight matrices and biases, $W_{ij}^{(l)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\frac{\sigma_{w}^{2}}{d_{l}})$ and $b_{i}^{(l)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma_{b}^{2})$ with $W_{ij}^{(l)}$ and $b_{i}^{(l)}$ being the matrix or vector entries of $\boldsymbol{W}^{(l)}$ and $\boldsymbol{b}^{(l)}$, respectively. As a shift operator, we choose
@@
where $\mathcal{A}$ is the adjacency matrix, $\mathbb{I}$ the identity in $\mathbb{R}^{N\times N}$, $D_{\alpha\beta}=\delta_{\alpha\beta}\sum_{\gamma}\mathcal{A_{\alpha\gamma}}$ is the degree matrix and $d_{\mathrm{max}}$ is the maximal degree. The parameter $g\in(0,1)$ allows us to weigh the off-diagonal elements compared to the diagonal ones. By construction the shift operator is column stochastic, which means that it has constant sums over columns $\sum_{\beta}A_{\alpha\beta}=\mathrm{const.}$. We will make use of this property in our analysis in Section 4.2. The generalization to non-stochastic shift operators is addressed in the discussion.
@@ 3.2 Gaussian process equivalence of GCNs
For simplicity we use $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ for which Equation (8) can be evaluated analytically; see Appendix B for details. It follows from (3) that
@@
In a semi-supervised node classification setting, we split the underlying graph into $N^{\mathrm{test}}$ unlabeled test nodes and $N^{\mathrm{train}}$ labeled training nodes ($N^{\mathrm{test}}+N^{\mathrm{train}}=N$); we correspondingly split the output random variable $Y$ into $Y_{\star}\in\mathbb{R}^{N^{\mathrm{test}}}$ and $Y_{D}\in\mathbb{R}^{N^{\mathrm{train}}}$. Features on the test nodes are predicted by conditioning on the values of the training nodes: $p\bigl{(}Y_{\ast}=y_{\ast}\mid Y_{D}=y_{D}\bigr{)}\,.$ This leads to the following posterior for the unobserved labels (see [31, 18] for details):
$Y_{\star}$ $\sim\mathcal{N}(m^{\mathrm{GP}},K^{\mathrm{GP}})\,,$ (10)
$m^{\mathrm{GP}}$ $=K_{\star D}^{(L+1)}(K_{DD}^{(L+1)}+\mathbb{I}\sigma_{ro}^{2})^{-1}Y_{D}\,,$ (11)
@@ 4.1 Propagation depths
We are interested in analyzing GCNs at large depth. We a priori assume that at infinite depth the GCN converges to an equilibrium in which covariances are static over layers $K_{\alpha\beta}^{(l)}\xrightarrow{l\to\infty}K^{\mathrm{eq}}$, irrespective of whether the GCN is in the oversmoothing or the chaotic phase. A posteriori we show that this assumption indeed holds. Since the fixed point $K^{\mathrm{eq}}$ is independent of the input, a GCN at equilibrium cannot use information from the input to make predictions (although, as we will see, in the non-oversmoothing phase it can still use the graph structure). In the following we analyze the equilibrium covariance $K^{\mathrm{eq}}$ to which GCNs with different $\sigma_{w}^{2}$, $\sigma_{b}^{2}$ and $A$ converge to, how they behave near this equilibrium, and at which rate it is approached.
@@
$\Delta_{\alpha\beta}^{(l+1)}$ $=\sum_{\gamma\delta}H_{\alpha\beta,\gamma\delta}\Delta_{\gamma\delta}^{(l)}+\mathcal{O}((\Delta^{(l)})^{2})\,,$ (16)
$H_{\alpha\beta,\gamma\delta}$ $=\sigma_{w}^{2}\sum_{\theta\phi}\frac{1}{2}(1+\delta_{\gamma,\delta})A_{\alpha\theta}A_{\beta\phi}\frac{\partial C_{\theta\phi}}{\partial K_{\gamma\delta}}[K^{\mathrm{eq}}]\,,$ (17)
where we use square brackets to denote the point around which we linearize.
@@
thus decoupling the system. For details on the linearization and some properties of the transition matrix $H$ refer to Appendix A.
@@ 4.2 The non-oversmoothing phase of GCNs
In this section we establish the chaotic, non-oversmoothing phase of GCNs, and show that this phase can be reached by simple tuning of the weight variance $\sigma_{w}^{2}$ at initialization. We start by noticing that a GCN is at a state of zero feature distance $\mu(\boldsymbol{X}^{(l)})=0$, if the covariance matrix has constant entries, $K_{\alpha\beta}^{(l)}=k^{(l)}$: Constant entries in $K_{\alpha\beta}^{(l)}$ imply that all preactivations are the same, $\boldsymbol{h}_{\alpha}^{(l)}=\boldsymbol{h}_{\beta}^{(l)}$, which in turn implies $C_{\alpha\beta}^{(l)}=c^{(l)}$ (by Equation (8)); the latter is equivalent to features being the same, $\boldsymbol{x}_{\alpha}^{(l)}=\boldsymbol{x}_{\beta}^{(l)}$. Due to our choice of the shift operator, the state of zero distance (and thus of $K_{\alpha\beta}^{(l)}=k^{(l)}$) is a always fixed point. Assuming that $C_{\alpha\beta}^{(l)}=c^{(l)}$, we obtain
@@ 4.2.1 Complete graph
Figure 1 shows how a GCN on a complete graph can be engineered to be non-oversmoothing by simple tuning of the weight variance $\sigma_{w}^{2}$. Panel a) shows the feature distance $\mu(\boldsymbol{X})$ in equilibrium as a heatmap dependent on the weight variance $\sigma_{w}^{2}$ and the parameter $g$ which determines the size of the off-diagonal elements of the shift operator $A$ (6). The red line is the theoretical prediction for the transition to the non-oversmoothing phase as calculated in Appendix C. The distance $\mu(\boldsymbol{X})$ increases from zero to finite values away from the transition into the non-oversmoothing phase. For larger values of $g$, the transition to chaos happens at larger values of $\sigma_{w}^{2}$: This makes sense intuitively, because smaller values of $g$ increase smoothing, so higher values of $\sigma_{w}^{2}$ are needed to counteract. The qualitative difference between the two phases is emphasized in panel b), where we depict $\mu(\boldsymbol{X})$ in equilibrium close to zero as black and white else. Thus our formalism correctly predicts the transition to the non-oversmoothing phase, allowing us to control how strongly the GCN oversmoothes. Panel c) shows the evolution of feature distance (13) over layers for randomly generated input features for different values of $\sigma_{w}^{2}$ and $g$ in a finite-size GCN and compares it to the GP prediction. The network parameters are marked with crosses of matching colors in panel b). While the network converges to zero feature distance in the oversmoothing phase, it converges to finite values of $\mu(\boldsymbol{X})$ in the chaotic, non-oversmoothing phase, meaning that in these cases the network does not oversmooth.
Figure 1: Simulations and GP prior of a GCN on a complete graph with $N=5$ nodes, shift operator $A_{\alpha\beta}=\frac{g}{N-1}+\delta_{\alpha\beta}(1-\frac{Ng}{N-1})$, vanishing bias $\sigma_{b}^{2}=0$ and $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$. a) The phase diagram dependent on $\sigma_{w}^{2}$ and $g$. The equilibrium feature distance $\mu(\boldsymbol{X})$ obtained from computing the GCN GP prior for $L=4,000$ layers is shown as a heatmap, the red line is the theoretical prediction for the transition to the non-oversmoothing phase. b) Same as in a) but color coding shows whether $\mu(\boldsymbol{X})$ is close to zero (black) or not (white) with precision $10^{-5}$. The red line again shows the theoretically predicted phase transition. c) Feature distance $\mu(\boldsymbol{X}^{(l)})$ for a random input $X_{\alpha i}^{(0)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1)$ as a function of layer $l$ for $(\sigma_{w}^{2},g)=(2,0.18),(3,0.14)$ (green) and $(\sigma_{w}^{2},g)=(4,0.1),(5,0.06),(6,0.02)$ (purple). The parameters are marked with color coded crosses in the phase diagram in panel b). Feature dimension of the hidden layers is $d_{l}=200$, crosses show the mean of $50$ network realizations, solid curves the theoretical predictions.

### v2

# Graph Neural Networks Do Not Always Oversmooth

@@ Abstract
Graph neural networks (GNNs) have emerged as powerful tools for processing relational data in applications. However, GNNs suffer from the problem of oversmoothing, the property that features of all nodes exponentially converge to the same vector over layers, prohibiting the design of deep GNNs. In this work we study oversmoothing in graph convolutional networks (GCNs) by using their Gaussian process (GP) equivalence in the limit of infinitely many hidden features. By generalizing methods from conventional deep neural networks (DNNs), we can describe the distribution of features at the output layer of deep GCNs in terms of a GP: as expected, we find that typical parameter choices from the literature lead to oversmoothing. The theory, however, allows us to identify a new, non-oversmoothing phase: if the initial weights of the network have sufficiently large variance, GCNs do not oversmooth, and node features remain informative even at large depth. We demonstrate the validity of this prediction in finite-size GCNs by training a linear classifier on their output. Moreover, using the linearization of the GCN GP, we generalize the concept of propagation depth of information from DNNs to GCNs. This propagation depth diverges at the transition between the oversmoothing and non-oversmoothing phase. We test the predictions of our approach and find good agreement with finite-size GCNs. Initializing GCNs near the transition to the non-oversmoothing phase, we obtain networks which are both deep and expressive.
@@ 1 Introduction
This approach allows us to extend the concept of a regular and a chaotic phase to GCNs: in the regular phase, which describes most of the GCNs studied in the current literature, distances between node features shrink over layers and exponentially attain the same value. We therefore call this the oversmoothing phase. On the other hand, if one increases the variance of the weights at initialization, it is possible to transition into the chaotic phase. In this phase, distances between node features converge to a fixed but finite distance at infinite depth. The convergence point is fully determined by the underlying graph structure and the hyperparameters of the GCN and may differ for different pairs of nodes. GCNs initialized in this phase thus do not suffer from oversmoothing. We find that the convergence point is informative about the topology of the underlying graph and may be used for node classification with GCNs of more than $1,000$ layers. Near the transition point, GCNs at large depth offer a trade-off between feature information and information contained in the neighborhood relation of the graph. We test the predictions of this theory and find good agreement in comparison to finite-size GCNs applied to the contextual stochastic block model [8]. On the citation network Cora [19] we reach the performance reported in the original work by Kipf and Welling [19] beyond $100$ layers. Our approach applies to graphs with arbitrary topologies, depths, and non-linearities.
@@ 2 Related Work
Oversmoothing is a well-known challenge within the GNN literature [21, 35]. On the theoretical side, rigorous techniques have been used to prove that oversmoothing is inevitable. The authors in [30] show that GCNs with the ReLU non-linearity exponentially lose expressive power and only carry information of the node degree in the infinite layer limit. While the authors notice that their upper bound for oversmoothing does not hold for large singular values of the weight matrices, they do not identify a non-oversmoothing phase in their model. These results have been extended in [2] to handle non-linearities different from ReLU. Also graph attention networks have been proven to oversmooth inevitably [45]. Their proof, however, makes assumptions on the weight matrices which, as we will show, exclude networks in the chaotic, non-oversmoothing phase.
@@ 3.1 Network architecture
with $\phi$ an elementwise non-linear activation function, $\boldsymbol{b}^{(l)}$ an optional bias term, weight matrices $\boldsymbol{W}^{(l)}\in\mathbb{R}^{d_{l}\times d_{l-1}}$, and $1\in\mathbb{R}^{N}$ a vector of all ones. We note that many GNN architectures studied in the literature are unbiased, which can be recovered by setting $\boldsymbol{b}^{(l)}$ to zero. We use a noisy linear readout, so that the output of the network is given by
$\boldsymbol{Y}=A\boldsymbol{X}^{(L)}\boldsymbol{W}^{(L+1)\top}+1\boldsymbol{b}^{(L+1)\top}+\boldsymbol{\epsilon}\,,$ (2)
with $\boldsymbol{\epsilon}\in\mathbb{R}^{N\times d_{L+1}}$ being independent Gaussian random variables: $\epsilon_{\alpha,i}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma_{ro}^{2})$. The readout noise $\boldsymbol{\epsilon}$ is included both to promote robust outputs and to prevent numerical issues in the matrix inversion in Equations (11) and (12). We use $d_{L+1}$ to denote the dimension of outputs.
@@
$\boldsymbol{y}_{\alpha}$ $=\boldsymbol{h}_{\alpha}^{(L+1)}+\boldsymbol{\epsilon}_{\alpha}\,,$ (5)
where $\boldsymbol{x}_{\alpha}^{(l)}\in\mathbb{R}^{d_{l}}$ is the feature vector of node $\alpha$ in layer $l$ and $\boldsymbol{y}_{\alpha}\in\mathbb{R}^{d_{L+1}}$ the network output for node $\alpha$. The values $\boldsymbol{h}_{\alpha}^{(l)}\in\mathbb{R}^{d_{l}}$ are linear functions of the features $\boldsymbol{x}_{\beta}^{(l-1)}$ and represent the input to the activation functions; we therefore refer to them as preactivations. The non-linearity $\phi(x)$ is applied elementwise to the preactivations $\boldsymbol{h}_{\alpha}^{(l)}$. While we leave $\phi(x)$ general for the development of the theory, we use $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ for the experiments in Section 4; this choice allows us to carry out certain integrals analytically. The scaling factor in the $\mathrm{erf}$ is chosen such that $\frac{\partial\phi}{\partial x}(0)=1$. We use independent and identical Gaussian priors for all weight matrices and biases, $W_{ij}^{(l)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\frac{\sigma_{w}^{2}}{d_{l}})$ and $b_{i}^{(l)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma_{b}^{2})$ with $W_{ij}^{(l)}$ and $b_{i}^{(l)}$ being the matrix or vector entries of $\boldsymbol{W}^{(l)}$ and $\boldsymbol{b}^{(l)}$, respectively. As a shift operator, we choose
@@
where $\mathcal{A}$ is the adjacency matrix, $\mathbb{I}$ the identity in $\mathbb{R}^{N\times N}$, $D_{\alpha\beta}=\delta_{\alpha\beta}\sum_{\gamma}\mathcal{A_{\alpha\gamma}}$ is the degree matrix and $d_{\mathrm{max}}$ is the maximal degree. The parameter $g\in(0,1)$ allows us to weigh the off-diagonal elements compared to the diagonal ones. By construction the shift operator is row-stochastic, which means that it has constant sums over columns $\sum_{\beta}A_{\alpha\beta}=1$. We will make use of this property in our analysis in Section 4.2. The generalization to non-stochastic shift operators will be shortly addressed later.
@@ 3.2 Gaussian process equivalence of GCNs
For simplicity we use $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ for which Equation (8) can be evaluated analytically; see Appendix A for details. It follows from (3) that
@@
In a semi-supervised node classification setting, we split the underlying graph into $N^{\mathrm{test}}$ unlabeled test nodes and $N^{\mathrm{train}}$ labeled training nodes ($N^{\mathrm{test}}+N^{\mathrm{train}}=N$); we correspondingly split the output random variable $Y_{i}\in\mathbb{R}^{N}$ for output dimension $i$ into $Y_{i}^{\star}\in\mathbb{R}^{N^{\mathrm{test}}}$ and $Y_{i}^{D}\in\mathbb{R}^{N^{\mathrm{train}}}$. Features on the test nodes are predicted by conditioning on the values of the training nodes: $p\bigl{(}Y_{i}^{\ast}=y_{i}^{\ast}\mid Y_{i}^{D}=y_{i}^{D}\bigr{)}\,.$ This leads to the following posterior for the unobserved labels (see [34, 20] for details):
$Y_{i}^{\star}$ $\sim\mathcal{N}(m_{i}^{\mathrm{GP}},K^{\mathrm{GP}})\,,$ (10)
$m_{i}^{\mathrm{GP}}$ $=K_{\star D}^{(L+1)}(K_{DD}^{(L+1)}+\mathbb{I}\sigma_{ro}^{2})^{-1}Y_{i}^{D}\,,$ (11)
@@ 4.1 Propagation depths
We are interested in analyzing GCNs at large depth. We a priori assume that at infinite depth the GCN converges to an equilibrium in which covariances are static over layers $K_{\alpha\beta}^{(l)}\xrightarrow{l\to\infty}K_{\alpha\beta}^{\mathrm{eq}}$, irrespective of whether the GCN is in the oversmoothing or the chaotic phase. A posteriori we show that this assumption indeed holds. Since the fixed point $K^{\mathrm{eq}}$ is independent of the input, a GCN at equilibrium cannot use information from the input to make predictions (although, as we will see, in the non-oversmoothing phase it can still use the graph structure). In the following we analyze the equilibrium covariance $K^{\mathrm{eq}}$ to which GCNs with different $\sigma_{w}^{2}$, $\sigma_{b}^{2}$ and $A$ converge to, how they behave near this equilibrium, and at which rate it is approached.
@@
$\Delta_{\alpha\beta}^{(l+1)}$ $=\sum_{\gamma,\delta}H_{\alpha\beta,\gamma\delta}\Delta_{\gamma\delta}^{(l)}+\mathcal{O}((\Delta^{(l)})^{2})\,,$ (16)
$H_{\alpha\beta,\gamma\delta}$ $=\sigma_{w}^{2}\sum_{\theta,\phi}\frac{1}{2}(1+\delta_{\gamma,\delta})A_{\alpha\theta}A_{\beta\phi}\frac{\partial C_{\theta\phi}}{\partial K_{\gamma\delta}}[K^{\mathrm{eq}}]\,,$ (17)
where we use square brackets to denote the point around which we linearize. The factor $\frac{1}{2}(1+\delta_{\gamma,\delta})$ is introduced to correctly count on- and off-diagonal elements of the covariance matrix, while the shift operators $A$ and the derivative $\frac{\partial C_{\theta\phi}}{\partial K_{\gamma\delta}}[K^{\mathrm{eq}}]$ originate from the message passing and the non-linearity $\phi$, respectively. The latter would result in a Kronecker delta $\frac{\partial C_{\theta\phi}}{\partial K_{\gamma\delta}}[K^{\mathrm{eq}}]=\delta_{\theta\phi,\gamma\delta}$ for linear networks. The calculation for $H_{\alpha\beta,\gamma\delta}$ is done in detail in Appendix B.
@@
thus decoupling the system. For details on the linearization and some properties of the transition matrix $H$ refer to Appendix B.
@@ 4.2 The non-oversmoothing phase of GCNs
In this section we establish the chaotic, non-oversmoothing phase of GCNs, and show that this phase can be reached by simple tuning of the weight variance $\sigma_{w}^{2}$ at initialization. We start by noticing that a GCN is at a state of zero feature distance $\mu(\boldsymbol{X}^{(l)})=0$, if the covariance matrix has constant entries, $K_{\alpha\beta}^{(l)}=k^{(l)}$: Constant entries in $K_{\alpha\beta}^{(l)}$ imply that all preactivations are the same, $\boldsymbol{h}_{\alpha}^{(l)}=\boldsymbol{h}_{\beta}^{(l)}$, which in turn implies $C_{\alpha\beta}^{(l)}=c^{(l)}$ (by Equation (8)); the latter is equivalent to features being the same, $\boldsymbol{x}_{\alpha}^{(l)}=\boldsymbol{x}_{\beta}^{(l)}$. Due to our choice of the shift operator, the state of zero distance (and thus of $K_{\alpha\beta}^{(l)}=k^{(l)}$) is always a fixed point. Assuming that $C_{\alpha\beta}^{(l)}=c^{(l)}$, we obtain
@@ 4.2.1 Complete graph
Figure 1: Simulations and GP prior of a GCN on a complete graph with $N=5$ nodes, shift operator $A_{\alpha\beta}=\frac{g}{N-1}+\delta_{\alpha\beta}(1-\frac{Ng}{N-1})$, vanishing bias $\sigma_{b}^{2}=0$ and $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$. a) The phase diagram dependent on $\sigma_{w}^{2}$ and $g$. The equilibrium feature distance $\mu(\boldsymbol{X})$ obtained from computing the GCN GP prior for $L=4,000$ layers is shown as a heatmap, the red line is the theoretical prediction for the transition to the non-oversmoothing phase. b) Same as in a) but color coding shows whether $\mu(\boldsymbol{X})$ is close to zero (black) or not (white) with precision $10^{-5}$. The red line again shows the theoretically predicted phase transition. c) Feature distance $\mu(\boldsymbol{X}^{(l)})$ for a random input $X_{\alpha i}^{(0)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,1)$ as a function of layer $l$. Parameters are written in the panel in matching colors and marked with color coded crosses in the phase diagram in panel b). Feature dimension of the hidden layers is $d_{l}=200$, crosses show the mean of $50$ network realizations, solid curves the theoretical predictions.
@@
Figure 1 shows how a GCN on a complete graph can be engineered to be non-oversmoothing by simple tuning of the weight variance $\sigma_{w}^{2}$. Panel a) shows that increasing the weight variance $\sigma_{w}^{2}$ or decreasing the size of the off-diagonal elements $g$ both shift the network towards the non-oversmoothing phase. Both parameters also increase the equilibrium feature distance beyond the transition. The theoretical prediction for the transition is calculated in Appendix C and shown as the red line. Panel b) confirms the accuracy of this calculation. Larger values of $g$ increase smoothing, and thus larger values of $\sigma_{w}^{2}$ are needed to compensate. Moreover, our formalism allows us to predict the evolution of feature distances over layers correctly, as can be confirmed in panel c). We find again that GCNs with parameters past the transition do not oversmooth.